Find the exact values by applying the odd/even identities, cofunction identities, and cosine of sum or difference identities. Do not use a calculator.
sin(85°)sin(40°)+sine(-5°)sin(-50°)

Answer by lwsshak3(11628)

Find the exact values by applying the odd/even identities, cofunction identities, and cosine of sum or difference identities. Do not use a calculator.
sin(85°)sin(40°)+sine(-5°)sin(-50°)
***
sin85º=cos(90-85)=cos5º
sin40º=cos(90-40)=cos50º
sin(-5º)=-sin5º
sin(-50º)=-sin50º
Identity: cos(s-t)=cos s cos t+sin s sin t
cos(50-5)=cos50º*cos5º+(-sin50º)*(-sin5º)
cos(50-5)=cos45º=√2/2
..
calculator check:
cos50º*cos5º+(-sin50º)*(-sin5º)≈0.6403+0.0667≈0.7071
Exact value=√2/2≈0.7071